A Runway 84 Restaurant reported the following current assets: cash $12,000.00, credit card receivables $1,800.00, accounts receivable $180.00, food inventory $4,400.00, and prepaid expenses, $1,120.00. Total current liabilities are $7,800.00. Calculate its current and acid-test ratios RESPECTIVELY.
Answer #1

Current Ratio = Current assets/Current liabilities

= (12,000+1,800+180+4400+1120)/7800

= 19,500/7800

= 2.5 times

Acid test ratio = Quick assets/current liabilities

= (19,500-4,400)/7,800

= 1.94 times